OpenStudy (anonymous):

A sum of $120,000 now at an interest rate of 10% per year compounded semiannually is equivalent to how much money 6 years ago? Solve using the effective annual ate and the effective semiannual rate.

OpenStudy (wolf1728):

So, basically you are asking what was the value of $120,000 (10% comp semi-annual) six yeas ago? Comp interest formula is: Total = Principal × ( 1 + Rate )^years we can rephrase that as: Principal = Total ÷ (1 + Rate)^years 10% compounded annually = 10.25% semi-annually Principal = 120,000 ÷ (1+ .01025)^6 Principal = 120,000 ÷ 1.0630976416 Principal = 112,877.68 six years ago To check that, let's see that amount compounded for 6 years. Total = Principal × ( 1 + Rate )^years Total = 112,877.68 x (1.01025)^6 Total = 120,000